OpenStudy (anonymous):

Explain Bohr's and Bury's rules for distribution of electrons into different shells

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@Siddharth18 @Preetha @SinginDaCalc2Blues

OpenStudy (anonymous):

@yrelhan4 @uri @geerky42 @AccessDenied

OpenStudy (anonymous):

pls help

OpenStudy (yrelhan4):

what does your textbook say?

OpenStudy (anonymous):

nothing! man I am just trying it but damn the ... leave it i dont know

OpenStudy (chmvijay):

it is also called n+l rule

OpenStudy (chmvijay):

According to this principle the orbital with the lowest energy will be filled first. The orbital having lower (n + l) value has lower energy. However for orbitals whose (n + l) values are equal, the orbital having lower value of 'n' has lower energy.

OpenStudy (yrelhan4):

thats one of the 3 rules.

OpenStudy (chmvijay):

what are other two

OpenStudy (yrelhan4):

hunds multiplicity and pauli's exclusion.

OpenStudy (chmvijay):

for example 3d and 4 s n(3)+L(2) =5 for 3d and for 4s =n(4) +l(1) =5

OpenStudy (chmvijay):

u left another one

OpenStudy (chmvijay):

aufbau principle

OpenStudy (chmvijay):

the question was to explain only bohrs and burys law

OpenStudy (yrelhan4):

n+l rule and aufbau are the same. and i dont know if bohrs and bury law involves only aufbau(n+l).

OpenStudy (chmvijay):

auf bau and n+L are not the same they have two different streams

OpenStudy (yrelhan4):

the order in which the orbitals are filled are filled is given by n+l rule right? what does aufbau give?

OpenStudy (chmvijay):

yaaa its one part of it ur right